# AnimeDL-2M: Million-Scale AI-Generated Anime Image Detection and Localization in Diffusion Era

### v1

@@ 5.1. Benchmark Settings
Protocol-CAT MVSS (Chen et al., 2021) 0.0295 0.0169 0.6423 0.5231
Protocol-CAT PSCC (Liu et al., 2022a) 0.0831 0.0553 0.5295 0.3601
@@
Protocol-CAT TruFor (Guillaro et al., 2023) 0.0235 0.0161 0.8470 0.7411
Protocol-CAT MMFusion (Triaridis and Mezaris, 2024) 0.1106 0.0771 1.0000 1.0000
@@
Table 3. Zero-shot results on AnimeDL-2M. GRE refers to (Sun et al., 2024a).
@@
MVSS (Chen et al., 2021) 0.8802 0.8423 1.0000 1.0000
PSCC (Liu et al., 2022a) 0.9398 0.9106 0.9994 0.9989
@@
TruFor (Guillaro et al., 2023) 0.9614 0.9362 0.9962 0.9925
MMFusion (Triaridis and Mezaris, 2024) 0.8615 0.8247 1.0000 1.0000
@@
AniXplore 0.9710 0.9506 1.0000 1.0000
AniXplore (HR) 0.9761 0.9923 1.0000 1.0000
Table 4. Comparison of our model and existing SOTA IMDL models on AnimeDL-2M. Our model is trained from scratch. ”HR” denotes high resolution version. For baseline models, since their scores when trained from scratch are either identical to finetuning results or less than 0.1% lower, we only report results finetuning results.
@@ 5.2. Implementation Details
Task Method Seen (FLUX) Unseen (SD) Unseen (SDXL) Seen (SD) Unseen (SDXL) Unseen (FLUX)
Localization MVSS 0.9264 / 0.8893 0.0026 / 0.0015 0.0164 / 0.0129 0.9298 / 0.8953 0.5581 / 0.5135 0.0852 / 0.0577
PSCC 0.9254 / 0.8883 0.0009 / 0.0006 0.0031 / 0.0022 0.9048 / 0.8684 0.0167 / 0.0108 0.0120 / 0.0074
Mesorch_P 0.9616 / 0.9362 0.0006 / 0.0005 0.0366 / 0.0339 0.9609 / 0.9372 0.0644 / 0.0512 0.0210 / 0.0133
AniXplore 0.9625 / 0.9375 0.0002 / 0.0001 0.0197 / 0.0175 0.8940 / 0.8535 0.1767 / 0.1584 0.0315 / 0.0206
AniXplore (HR) 0.9774 / 0.9603 0.0000 / 0.0000 0.0005 / 0.0004 0.9244 / 0.8841 0.1722 / 0.1198 0.0329 / 0.0213
Detection MVSS 0.5460 / 0.3755 0.1139 / 0.0604 0.1204 / 0.0641 0.9984 / 0.9967 0.9442 / 0.8944 0.8811 / 0.7875
PSCC 0.5038 / 0.3367 0.0082 / 0.0041 0.0101 / 0.0051 0.9984 / 0.9969 0.6824 / 0.5179 0.6702 / 0.5039
AniXplore 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000
AniXplore (HR) 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000 1.0000 / 1.0000
Table 5. Cross-dataset evaluation results. Left columns show performance when trained on FLUX; right columns show performance when trained on SD. Metrics are F1 / IoU for localization and F1 / Accuracy for detection. HR denotes high resolution version.
@@ 5.4. Cross-Dataset and In-the-Wild Evaluation
Factors Influencing Generalization. As shown in Table 5, all models exhibit generally poor performance on the localization task under cross-dataset settings. This is likely because localization heavily relies on identifying artifact patterns left by AI-generated regions. However, such artifacts vary significantly across generative models, making it difficult to generalize. This also suggests that training or fine-tuning on the target domain can substantially improve localization performance, which is consistent with the findings in (Epstein et al., 2023). Combining results in Table 6, we can see that generalize ability on detection task does not strongly correlate with localization performance. Some models achieve high detection accuracy across domains despite limited localization ability. This implies that detection generalization may depend more on the robustness of the model architecture than on the ability to capture specific forgery artifacts. Among all tested methods, PSCC (Liu et al., 2022a), as the only model that uses RGB images as the sole input modality, demonstrates the weakest generalization, highlighting the importance of multi-modal or multi-channel feature inputs for generalizability. Meanwhile, both TruFor (Guillaro et al., 2023) and MMFusion (Triaridis and Mezaris, 2024) incorporate noise-based features, yet their performance differs significantly. This suggests that not all handcrafted features are equally effective, and the design of the feature extractor plays a critical role in mitigating overfitting. Therefore, careful selection and design of input features is essential for building more generalizable forensic models.

### v2

@@ 5.1. Benchmark Settings
Protocol-CAT MVSS (Chen et al., 2021) 0.0295 0.0169 0.8627 0.7587
Protocol-CAT PSCC (Liu et al., 2022a) 0.0831 0.0553 0.4808 0.4008
@@
Protocol-CAT TruFor (Guillaro et al., 2023) 0.0235 0.0161 0.8467 0.7408
Protocol-CAT MMFusion (Triaridis and Mezaris, 2024) 0.1106 0.0771 0.6941 0.5784
@@
Table 3. Zero-shot results on AnimeDL-2M. ”Pretrain” denotes the dataset used for pretraining. GRE refers to (Sun et al., 2024a).
@@
MVSS (Chen et al., 2021) 0.8802 0.8423 0.8991 0.8269
PSCC (Liu et al., 2022a) 0.9398 0.9106 0.9993 0.9989
@@
TruFor (Guillaro et al., 2023) 0.9614 0.9362 0.9959 0.9938
MMFusion (Triaridis and Mezaris, 2024) 0.8615 0.8247 0.9581 0.9463
@@
AniXplore 0.9710 0.9506 0.9999 0.9998
AniXplore (HR) 0.9761 0.9923 0.9999 0.9999
Table 4. Comparison of our model and existing SOTA IMDL models on AnimeDL-2M. Our model is trained from scratch. ”HR” denotes high resolution version. For baseline models, we presents results that finetuned from Protocol-CAT checkpoint. Since pixel-level F1 score will be 0 for real images, pixel-level test only includes fake images while image-level test contains both real and fake images.
@@ 5.2. Implementation Details
Task Method Seen (FLUX) Unseen (SD) Unseen (SDXL) Real Seen (SD) Unseen (SDXL) Unseen (FLUX) Real
Loc. MVSS 0.9264 / 0.8893 0.0026 / 0.0015 0.0164 / 0.0129 / 0.9298 / 0.8953 0.5581 / 0.5135 0.0852 / 0.0577 /
PSCC 0.9254 / 0.8883 0.0009 / 0.0006 0.0031 / 0.0022 / 0.9048 / 0.8684 0.0167 / 0.0108 0.0120 / 0.0074 /
Mesorch_P 0.9616 / 0.9362 0.0006 / 0.0005 0.0366 / 0.0339 / 0.9609 / 0.9372 0.0644 / 0.0512 0.0210 / 0.0133 /
TruFor 0.9621 / 0.9365 0.0015 / 0.0011 0.0520 / 0.0459 / 0.9617 / 0.9379 0.0627 / 0.0468 0.0388 / 0.0244 /
AniXplore 0.9625 / 0.9375 0.0002 / 0.0001 0.0197 / 0.0175 / 0.8940 / 0.8535 0.1767 / 0.1584 0.0315 / 0.0206 /
AniXplore (HR) 0.9774 / 0.9603 0.0000 / 0.0000 0.0005 / 0.0004 / 0.9244 / 0.8841 0.1722 / 0.1198 0.0329 / 0.0213 /
Det. MVSS 0.5460 / 0.3755 0.1139 / 0.0604 0.1204 / 0.0641 0.9143 0.9984 / 0.9967 0.9442 / 0.8944 0.8811 / 0.7875 0.9950
PSCC 0.5038 / 0.3367 0.0082 / 0.0041 0.0101 / 0.0051 0.9974 0.9984 / 0.9969 0.6824 / 0.5179 0.6702 / 0.5039 0.9963
TruFor 0.5468 / 0.3763 0.0206 / 0.0104 0.1211 / 0.0644 0.9997 0.9977 / 0.9953 0.7490 / 0.5988 0.6586 / 0.4909 0.9968
AniXplore 0.5107 / 0.3429 0.0021 / 0.0010 0.0283 / 0.0144 0.9997 0.9989 / 0.9979 0.7922 / 0.6559 0.7147 / 0.5560 0.9534
AniXplore (HR) 0.5002 / 0.3335 0.0001 / 0.0001 0.0004 / 0.0002 1.0000 0.9999 / 0.9997 0.8629 / 0.7588 0.7169 / 0.5587 0.9922
Table 5. Cross-dataset evaluation results. Left columns show performance when trained on FLUX; right columns show performance when trained on SD. Metrics are F1 / IoU for localization and F1 / Accuracy for detection, and Accuracy for ’real’ columns. HR denotes high resolution version.
@@ 5.4. Cross-Dataset and In-the-Wild Evaluation
Factors Influencing Generalization. As shown in Table LABEL:tab:cross-dataset-merged, all models exhibit generally poor performance on the localization task under cross-dataset settings. This is likely because localization heavily relies on identifying artifact patterns left by AI-generated regions. However, such artifacts vary significantly across generative models, making it difficult to generalize. This also suggests that training or fine-tuning on the target domain can substantially improve localization performance, which is consistent with the findings in (Epstein et al., 2023). Combining results in Table 6, we can see that generalize ability on detection task does not strongly correlate with localization performance. Some models achieve high detection accuracy across domains despite limited localization ability. This implies that detection generalization may depend more on the robustness of the model architecture than on the ability to capture specific forgery artifacts. Among all tested methods, PSCC (Liu et al., 2022a), as the only model that uses RGB images as the sole input modality, demonstrates the weakest generalization, highlighting the importance of multi-modal or multi-channel feature inputs for generalizability. Meanwhile, both TruFor (Guillaro et al., 2023) and MMFusion (Triaridis and Mezaris, 2024) incorporate noise-based features, yet their performance differs significantly. This suggests that not all handcrafted features are equally effective, and the design of the feature extractor plays a critical role in mitigating overfitting. Therefore, careful selection and design of input features is essential for building more generalizable forensic models.